A monoprotic weak acid, HA, dissociates in water according to the reaction

HA(aq)↽−−⇀H+(aq)+A−(aq)

The equilibrium concentrations of the reactants and products are [HA]=0.280 M, [H+]=2.00×10−4 M, and [A−]=2.00×10−4 M. Calculate the value of pKa for the acid HA.
Answer #1

Ka = [H+][A-]/[HA]
= (2.00*10^-4)*(2.00*10^-4)/(0.280)
= 1.43*10^-7


use:
pKa = -log Ka
= -log (1.43*10^-7)
= 6.84
Answer: 6.84